# Characterizing spin ordering via maximal row correlation in classical spin models

### v1

@@ I Introduction
The Potts model, introduced by R. B. Potts in his 1951 PhD thesis [3] and subsequent 1952 paper [4], is a fundamental model in statistical physics. It generalizes the Ising model by allowing each lattice site to occupy one of $q$ discrete states. Over the past half-century, it has been the subject of extensive study. For comprehensive reviews, see, e.g., Refs. 5, 6. The phase transition of the Potts model with ferromagnetic (FM) interaction is relatively straightforward, and its properties are well understood. In contrast, the behavior of the AFM Potts model is significantly more complex, with phase transition characteristics that depend on both the lattice structure and the number of states $q$. Some Potts models exhibit macroscopic ground-state degeneracy and zero-temperature entropy, potentially giving rise to phases where certain sites are ordered while others remain disordered. This phenomenon of partial order has been observed in both frustrated [7, 8, 9, 10] and non-frustrated models [11, 2], highlighting the difficulty of defining a clear order parameter for the AFM Potts model.
A common approach for studying partial orders is to compute the partial or staggered magnetization by adding or subtracting the magnetization of sublattices, which often requires model-specific analysis [12, 13]. For certain models, however, identifying the order from low-temperature configurations can be challenging, as seen in the case of the AFM Ising model on the Union-Jack lattice.
Recently, in contrast to conventional methods of identifying phase transitions and critical points through the study of thermodynamic quantities and order parameters, substantial progress has been made in exploring alternative approaches. These include the “pattern picture” [14], the perspective of information theory [15], analyzing configuration correlations in the Hilbert space during the Monte Carlo (MC) process [16, *Cheng_2024], and machine learning-based analyses [18, 19], among others. Inspired by these advancements, the present study aims to propose a unified and physically sound order parameter within a general framework to describe the phase transitions in various Potts models, particularly those with AFM interactions and no apparent definition of long-range orders.
@@ II.1 Maximal row correlation
Figure 1: (a) An example of the configuration matrix $A$ corresponding to a spin configuration of the Ising model on the $4\times 4$ square lattice, where the entries +1/1+1/-1+ 1 / - 1 represent the spin orientations (up and down, respectively) at the lattice sites. The maximal row correlation is calculated from the largest eigenvalue of the row correlation matrix $M=A\,A^{\mathrm{T}}$. (b) The purpose of the process is to linearly combine the rows of the matrix $A$ to extract the maximal degree of order present in the spin configuration.
@@
As illustrated in Fig. 1, any given spin configuration on the square lattice can be encoded in a matrix $A$ of the same dimensions, where each matrix elements records the spin orientation at the corresponding lattice site. We refer to this as the configuration matrix. In principle, any 2D lattice can be rearranged into a rectangular array, allowing the application of the configuration matrix representation. For higher-dimensional lattices, the corresponding configuration matrix can be constructed by reshaping higher-order tensors, for example by combining multiple indices into a single one. As a illustration, we refer the reader to Appendix B, where the 3D and 4D Potts models are discussed.
A positive semi-definite square matrix $M$ can be constructed from $A$ as $M=A\,A^{\mathrm{T}}$, with its dimension determined by the number of rows of $A$. The physical meaning of the matrix $M$ is clear: each entry $(M)_{ij}$ represents the inner product between the $i$-th and $j$-th rows of $A$, thereby quantifying their correlation. Therefore we call $M$ the row correlation matrix. The maximal row correlation, denoted as ${O}$ in the paper, is simply defined to be the square root of $\lambda_{\mathrm{M}}$ that is the largest eigenvalue of $M$, divided by the matrix size, i.e., $O=\sqrt{\lambda_{\mathrm{M}}}/L$ ($L$ refers to the number of rows of $A$), as illustrated in Fig. 1(a). Consequently, $\delta O$, which measures the fluctuation of $O$, is written as $\delta O=\left(\left\langle O^{2}\right\rangle-\left\langle O\right\rangle^{2}\right)\times L$. The reason for these rescalings is explained in Appendix A.
The matrix $M$, defined as the product of the configuration matrix $A$ with its transpose $A^{\mathrm{T}}$, is analogous to the reduced density matrix in the density-matrix renormalization group (DMRG) [20], apart from the normalization condition. In this analogy, the configuration matrix $A$ corresponds to the superblock wavefunction in the DMRG framework, with its row and column indices representing the basis states of the system and environment, respectively. In this sense, the eigenvalue distribution of $M$ may be interpreted as encoding the entanglement characteristics of the spin configurations between the row and column directions. As for the largest eigenvalue of $M$, a geometric understanding is following: let $\mathbf{v}$ be an eigenvector of $M$ with eigenvalue $\lambda$, i.e., $M\mathbf{v}=\lambda\mathbf{v}$, then we have $\mathbf{v}^{\mathrm{T}}A\,A^{\mathrm{T}}\mathbf{v}=\lambda\mathbf{v}^{\mathrm{T}}\mathbf{v}$. This indicates that the largest eigenvalue of $M$ corresponds to the maximal squared length among all vectors obtained by linear combinations of the row vectors of $A$, where the coefficients are normalized. It occurs when the coefficient vector $\mathbf{v}$ coincides with the eigenvector of $M$ associated with the largest eigenvalue $\lambda_{\mathrm{M}}$. This interpretation is illustrated in Fig. 1(b). That is why we refer to the order parameter $O$ as the maximal row correlation.
@@
By construction, the definition ensures that the average spin correlation between any two sites with random (i.e., uncorrelated) states is zero. When $q=2$, we can recover the familiar Ising case. Once $M$ is available, the order parameter $O$ can be calculated accordingly. As for the classical continuous spin model addressed in Appendix C, the component product between the spin configurations at two sites is defined as the inner (dot) product of their respective spin vectors.
@@ II.2 Monte Carlo simulation
In the numerical simulations, for the Potts model with AFM interactions, the standard classical MC method with the Metropolis algorithm [21] is employed. For each temperature, $200\times(q-1)\times L^{2}$ ($L$ is the system size) MC steps are performed for equilibration. (For the XY model and the Heisenberg model discussed in the Appendix C, $20{,}000\times L^{2}$ MC steps are used.) Each system size is simulated with $80$ independent annealing processes. In the afterward measurement phase, one sample is taken every $10\times L^{2}$ steps, yielding $1{,}000$ samples per process. This results in $80\times 1{,}000$ configurations per temperature, from which the order parameter $O$ is computed for each configuration. These results are then evenly divided into $80$ bins for statistical analysis. In the vicinity of the critical regime, MC simulations are conducted with finer sampling to improve numerical precision.
On the other hand, for the Potts models with FM interactions, the Wolff update [22] is implemented near the critical point to improve equilibration efficiency, allowing equilibration with only $L^{2}$ MC steps per temperature. In the measurement phase, one sample is taken every $10$ cluster updates.
In the presence of a phase transitions, analogous to the case of magnetization, the critical components—also denoted here as $\beta$ and $\nu$—along with the critical temperatures $T_{\mathrm{c}}$, are determined via a data collapse analysis of the maximal row correlation $O$. The procedure of data collapse is as follows. The data points for various lattice sizes near the critical regime are rescaled into a universal curve by choosing appropriate values of $T_{\mathrm{c}}$, $\beta$ and $\nu$ (for example, see Fig. 2(c) for the case of the $q=2$ FM Potts model on square lattice). The curve is fitted by a eleventh-order polynomial. We employ the binning method to estimate standard errors, and perform a chi-squared test [23] based on these errors to assess the quality of the data collapse and to optimize the parameters $T_{\mathrm{c}}$, $\beta$, $\nu$. Note that the chi-squared value is defined as
@@ III Results
Here $q$ is the number of the available states for each site. $\sigma_{i}$ denotes the state at the $i$-th site, commonly referred to as the spin. $\delta_{\sigma_{i},\sigma_{j}}$ is the Kronecker delta function, that takes the value of $1$ if $\sigma_{i}=\sigma_{j}$, and 00 otherwise. $J$ is the coupling constant between spins. When $J>0$, the coupling is FM, whereas $J<0$ corresponds to AFM coupling. In this study, for simplicity, only the couplings between the nearest neighbors are considered. When $q=2$, the Potts model simplifies to the Ising model. In the following discussion, the magnitude of the coupling constant $J$ is always set to unity.
III.1 $q=2$ and $q=3$ FM Potts model on square lattice
Figure 2: The temperature dependence of the maximal row correlation $O$ (a), and its fluctuations $\delta O$ (b) for the $q=2$ FM Potts model on square lattice with various sizes. The vertical dashed lines in (a) and (b) mark the position of the exact critical temperature $T_{c}$ in the thermodynamic limit. The results of the data collapse with the optimized critical temperature and exponents with fluctuation range, including the value of $\chi^{2}$, are shown in (c). In (b), error bars are shown for large lattice size ($L=128$ here). For all other cases, including the data for $O$ in panel (a), error bars are omitted due to their negligible magnitude, and the size of the symbols does not reflect the actual uncertainty.
Figure 3: Same quantities as in Fig. 2, but for the $q=3$ FM Potts model on square lattice.
To benchmark the effectiveness of the maximal row correlation $O$ as the order parameter in classical spin systems, we first perform the MC simulations on the $q=2$ and $q=3$ FM Potts models on square lattice. The exact critical temperature in the thermodynamic limit of the $q$-state Potts model on the square lattice with isotropic FM coupling is known to be $T_{c}=J/\ln(1+\sqrt{q})$, where $J$ is the coupling strength [5, 6]. The critical exponents, relevant for our present study, are $\beta=1/8$, $\nu=1$ for the prestigious Ising case ($q=2$), and $\beta=1/9$, $\nu=5/6$ for the $q=3$ case [5, 6]. Our numerical MC results are presented in Figs. 2 and 3, respectively.
@@ III.1 $q=2$ and $q=3$ FM Potts model on square lattice
The above observations suggest a close relationship in the FM Potts model between the maximal row correlation $O$ and the magnetization $M$, as well as between $\delta O$ and the susceptibility $\chi$. It might be not surprising, given that $O$ is proportional to the square root of the maximal eigenvalue of the row-row correlation matrix $M$, as discussed in the Methods section II. The correspondence between $O$ and $M$ (as well as between $\delta O$ and $\chi$) can be further reinforced by the calculations of the relevant critical exponents. As a demonstration, we apply the data collapse method (detailed in Section II) to analyze the MC data of $O$ around the phase transition point for various lattice sizes. This allow us to determine the optimal values of $T_{c}$, $\beta$ and $\nu$, along with their uncertainty estimates. The results are presented in Figs. 2(c) and 3(c).
We see that both for the $q=2$ and $3$ cases, the critical temperatures obtained numerically are in good agreement with the exact results. As for the critical exponents, when $q=2$ (2D Ising case), we have in our calculation $\nu=0.96$ and $\beta=0.122$, where the known exact values are $\nu=1$, $\beta=1/8=0.125$; when $q=3$, the numerical results are $\nu=0.81$ and $\beta=0.105$, where the exact values are known to be $\nu=5/6\approx 0.833$, $\beta=1/9\approx 0.111$. The deviations we suggest are mainly due to the finite size effect. In the FM cases, the accuracy of using $O$ as the order parameter is comparable to that of the conventional magnetization $M$ for the same lattice sizes.
III.2 $q=3$ AFM Potts model on diced lattice
@@
Figure 4: The diced lattice, as shown in (a), consists of a honeycomb sublattice (red circles, majority sites) and a triangular sublattice (blue circles, minority sites), where the latter is formed by sites located at the centers of the hexagons. Each blue circle is connected to six neighboring red circles, forming a bipartite structure. In order to map the spin configurations on the diced lattice into matrix form, an oblique rectilinear coordinate frame is used. In this coordinate frame, a blue circle appears after every two consecutive red circles along either the $x$- or $y$-direction. In (b) a snapshot of the spin configuration on $L=18$ diced lattice for the $q=3$ AFM Potts model is shown. It is taken from the MC sampling after the equilibration at the temperature $T=1.0\times 10^{-5}$. Three colors (red, blue and yellow) represent three spin values. A FM long-range order of one spin component (blue) is clearly observed on the triangular sublattice. In contrast, the spins on the honeycomb sublattice appear to randomly adopt the other two values (red and yellow).
@@
Figure 5: Same quantities as in Fig. 2, but for the $q=3$ AFM Potts model on diced lattice. The vertical dashed lines in (a) and (b) mark the position of the critical temperature $T_{c}$ obtained in Ref. [12].
@@
The critical temperature $T_{c}$ and the critical exponents $\nu$ and $\beta$, determined by the data collapse method, are shown in Fig. 5(c). The estimated critical temperature $T_{c}=0.507$ is in excellent agreement with the reference values. The critical exponents, $\nu=0.77$, with an uncertainty range of $[0.75,0.79]$ and $\beta=0.110$, are also consistent with those of the $q=3$ FM Potts model ($\nu=5/6\approx 0.833$, $\beta=1/9\approx 0.111$). These observations corroborate the conclusion in Ref. [12] that the critical behavior of the model belongs to the universality class of the $q=3$ Potts ferromagnet.

### v2

@@ I Introduction
The Potts model, introduced by R. B. Potts in his 1951 PhD thesis [3] and subsequent 1952 paper [4], is a fundamental model in statistical physics. It generalizes the Ising model by allowing each lattice site to occupy one of $q$ discrete states. Over the past half-century, it has been the subject of extensive study. For comprehensive reviews, see, e.g., Refs. [5] and  [6]. The phase transition of the Potts model with ferromagnetic (FM) interaction is relatively straightforward, and its properties are well understood. In contrast, the behavior of the AFM Potts model is significantly more complex, with phase transition characteristics that depend on both the lattice structure and the number of states $q$. Some Potts models exhibit macroscopic ground-state degeneracy and zero-temperature entropy, potentially giving rise to phases where certain sites are ordered while others remain disordered. This phenomenon of partial order has been observed in both frustrated [7, 8, 9, 10] and non-frustrated models [11, 2], highlighting the difficulty of defining a clear order parameter for the AFM Potts model.
A common approach for studying partial orders is to compute the partial or staggered magnetization by adding or subtracting the magnetization of sublattices, which often requires model-specific analysis [12, 13]. For certain models, however, identifying the order from low-temperature configurations can be challenging, as seen in the case of the $q=2$ AFM Ising model on the Union-Jack lattice (see also Fig. 6 in the main text).
Recently, in contrast to conventional methods of identifying phase transitions and critical points through the study of thermodynamic quantities and order parameters, substantial progress has been made in exploring alternative approaches. These include the “pattern picture” [14], the perspective of information theory [15], analyzing configuration correlations during the Monte Carlo (MC) process [16, 17, *Cheng_2024], and machine learning-based analyses [19, 20], among others. Inspired by these advancements, the present study aims to propose a unified and physically sound order parameter within a general framework to describe the phase transitions in various Potts models, particularly those with AFM interactions and no apparent definition of long-range orders.
@@ II.1 Maximal row correlation
Figure 1: (a) An example of the configuration matrix AA corresponding to a spin configuration of the Ising model on the $4\times 4$ square lattice, where the entries +1/1+1/-1 represent the spin orientations (up and down, respectively) at the lattice sites. The maximal row correlation is calculated from the largest eigenvalue of the row correlation matrix $M=A\,A^{\mathrm{T}}$. (b) The purpose of the process is to linearly combine the rows of the matrix AA to extract the maximal degree of order present in the spin configuration.
@@
As illustrated in Fig. 1, any given spin configuration on the square lattice can be encoded in a matrix $A$ of the same dimensions, where each matrix elements records the spin orientation at the corresponding lattice site. We refer to this as the configuration matrix. In principle, any 2D lattice can be rearranged into a rectangular array, allowing the application of the configuration matrix representation. For higher-dimensional lattices, the corresponding configuration matrix can be constructed by reshaping higher-order tensors, for example by combining multiple indices into a single one. As a illustration, we refer the reader to Appendix C, where the 3D and 4D Potts models are discussed.
A positive semi-definite square matrix $M$ can be constructed from $A$ as $M=A\,A^{\mathrm{T}}$, with its dimension determined by the number of rows of $A$. The physical meaning of the matrix $M$ is clear: each entry $(M)_{ij}$ represents the inner product between the $i$-th and $j$-th rows of $A$, thereby quantifying their correlation. Therefore we call $M$ the row correlation matrix. The maximal row correlation, denoted as O{O} in the paper, is simply defined to be the square root of $\lambda_{\mathrm{M}}$ that is the largest eigenvalue of $M$, divided by the matrix size, i.e., $O=\sqrt{\lambda_{\mathrm{M}}}/L$ ($L$ refers to the number of rows of $A$), as illustrated in Fig. 1(a). Consequently, $\delta O$, which measures the fluctuation of $O$, is written as $\delta O=\left(\left\langle O^{2}\right\rangle-\left\langle O\right\rangle^{2}\right)\times L$. The reason for these rescalings is explained in Appendix A.
The geometric significance of $\lambda_{\mathrm{M}}$, the largest eigenvalue of the row correlation matrix $M$, is as follows. Let $\mathbf{v}$ be an eigenvector of $M$ with eigenvalue $\lambda$, such that $M\mathbf{v}=\lambda\mathbf{v}$. Then, $\mathbf{v}^{\mathrm{T}}A\,A^{\mathrm{T}}\mathbf{v}=\lambda\mathbf{v}^{\mathrm{T}}\mathbf{v}$. This relation implies that $\lambda_{\mathrm{M}}$ gives the maximal squared length of any linear combination of the row vectors of $A$ with normalized coefficients, i.e., under the constraint $\mathbf{v}^{\mathrm{T}}\mathbf{v}=1$. The maximum is attained when $\mathbf{v}$ is the eigenvector of $M$ corresponding to $\lambda_{\mathrm{M}}$. This interpretation is illustrated in Fig. 1(b). That is why we refer to the order parameter $O$ as the maximal row correlation. It is worth noting that the row correlation matrix $M$, derived from an individual spin configuration, is distinct from the standard covariance matrix employed in principal component analysis (PCA) [21], which is typically defined over an ensemble of configurations [22, 16].
@@
By construction, the definition ensures that the average spin correlation between any two sites with random (i.e., uncorrelated) states is zero. When $q=2$, we can recover the familiar Ising case. Once $M$ is available, the order parameter $O$ can be calculated accordingly. As for the classical continuous spin model addressed in Appendix D, the component product between the spin configurations at two sites is defined as the inner (dot) product of their respective spin vectors.
@@ II.2 Monte Carlo simulation
In the numerical simulations, for the Potts model with AFM interactions, the standard classical MC method with the Metropolis algorithm [23] is employed. For each temperature, $200\times(q-1)\times L^{2}$ ($L$ is the system size) MC steps are performed for equilibration. In each MC step, a single-site spin-flip attempt is made. Each system size is simulated with $80$ independent annealing processes. In the afterward measurement phase, one sample is taken every $10\times L^{2}$ steps, yielding 1,0001{,}000 samples per process. This results in $80\times 1{,}000$ configurations per temperature, from which the order parameter $O$ is computed for each configuration. These results are then evenly divided into $80$ bins for statistical analysis. In the vicinity of the critical regime, MC simulations are conducted with finer sampling to improve numerical precision.
On the other hand, for the Potts models with FM interactions, the Wolff update [24] is implemented near the critical point to improve equilibration efficiency, allowing equilibration with only $L^{2}$ cluster updates per temperature. In the measurement phase, one sample is taken every $10$ cluster updates. As in the AFM case, $80$ independent annealing processes are performed at each temperature, with 1,0001{,}000 samples collected per process.
In the presence of a phase transitions, analogous to the case of magnetization, the critical components—also denoted here as $\beta$ and $\nu$—along with the critical temperatures $T_{\mathrm{c}}$, are determined via a data collapse analysis of the maximal row correlation $O$. The procedure of data collapse is as follows. The data points for various lattice sizes near the critical regime are rescaled into a universal curve by choosing appropriate values of $T_{\mathrm{c}}$, $\beta$ and $\nu$ (for example, see Fig. 2(c) for the case of the $q=2$ FM Potts model on the square lattice). The curve is fitted by a eleventh-order polynomial. We employ the binning method to estimate standard errors, and perform a chi-squared test [25] based on these errors to assess the quality of the data collapse and to optimize the parameters $T_{\mathrm{c}}$, $\beta$, $\nu$. Note that the chi-squared value is defined as
@@
To better access the quality of data fit relative to statistical uncertainty, in the following data collapse analysis we calculate the reduced chi-squared value as
$\chi^{2}_{\text{red}}=\frac{\chi^{2}}{\text{dof}}.$ (4)
Here the degrees of freedom are given by $\text{dof}=N-p$, where $N$ is the number of data points and $p$ the number of fitting parameters. A value of $\chi^{2}_{\text{red}}$ close to unity typically indicates a good-quality fit. In our analysis, we use five system sizes for a given model, with $30$ data points per size near the critical temperature, yielding a total of $N=150$ data points. With the fitting curve described by an eleventh-order polynomial ($p=12$), this results in $\text{dof}=150-12=138$.
@@ III Results
Here $q$ is the number of the available states for each site. $\sigma_{i}$ denotes the state at the $i$-th site, commonly referred to as the spin. $\delta_{\sigma_{i},\sigma_{j}}$ is the Kronecker delta function, that takes the value of $1$ if $\sigma_{i}=\sigma_{j}$, and 0 otherwise. $J$ is the coupling constant between spins. When $J>0$, the coupling is FM, whereas $J<0$ corresponds to AFM coupling. In this study, for simplicity, only the couplings between the nearest neighbors are considered. When $q=2$, the Potts model simplifies to the Ising model. In the following discussion, the magnitude of the coupling constant $J$ is always set to unity.
III.1 $q=2$ and $q=3$ FM Potts model on the square lattice
Figure 2: The temperature dependence of the maximal row correlation $O$ (a), and its fluctuations $\delta O$ (b) for the $q=2$ FM Potts model on the square lattice with various sizes. The vertical dashed lines in (a) and (b) mark the position of the exact critical temperature $T_{c}$ in the thermodynamic limit. The results of the data collapse with the optimized critical temperature and exponents with fluctuation range, including the value of the reduced chi-squared $\chi^{2}_{\text{red}}$, are shown in (c). In (b), error bars are shown for large lattice size ($L=128$ here). For all other cases, including the data for $O$ in panel (a), error bars are omitted due to their negligible magnitude, and the size of the symbols does not reflect the actual uncertainty.
Figure 3: Same quantities as in Fig. 2, but for the $q=3$ FM Potts model on the square lattice.
To benchmark the effectiveness of the maximal row correlation $O$ as the order parameter in classical spin systems, we first perform the MC simulations on the $q=2$ and $q=3$ FM Potts models on the square lattice. The exact critical temperature in the thermodynamic limit of the $q$-state Potts model on the square lattice with isotropic FM coupling is known to be $T_{c}=J/\ln(1+\sqrt{q})$, where $J$ is the coupling strength [5, 6]. The critical exponents, relevant for our present study, are $\beta=1/8$, $\nu=1$ for the prestigious Ising case ($q=2$), and $\beta=1/9$, $\nu=5/6$ for the $q=3$ case [5, 6]. Our numerical MC results are presented in Figs. 2 and 3, respectively.
@@ III.1 $q=2$ and $q=3$ FM Potts model on the square lattice
The above observations suggest a close relationship in the FM Potts model between the maximal row correlation $O$ and the magnetization $M$, as well as between $\delta O$ and the susceptibility $\chi$. It might be not surprising, given that $O$ is proportional to the square root of the maximal eigenvalue of the row-row correlation matrix $M$, as discussed in the Methods section II. The correspondence between $O$ and $M$ (as well as between $\delta O$ and $\chi$) can be further reinforced by the calculations of the relevant critical exponents. As a demonstration, we apply the data collapse method (detailed in Sec. II) to analyze the MC data of $O$ around the phase transition point for various lattice sizes. This allow us to determine the optimal values of $T_{c}$, $\beta$ and $\nu$, along with their uncertainty estimates. The results are presented in Figs. 2(c) and 3(c).
We see that both for the $q=2$ and $3$ cases, the critical temperatures obtained numerically are in good agreement with the exact results. As for the critical exponents, when $q=2$ (2D Ising case), we have in our calculation $\nu=0.96$ and $\beta=0.122$, where the known exact values are $\nu=1$, $\beta=1/8=0.125$; when $q=3$, the numerical results are $\nu=0.81$ and $\beta=0.105$, where the exact values are known to be $\nu=5/6\approx 0.833$, $\beta=1/9\approx 0.111$. The deviations we suggest are mainly due to the finite size effect. In the FM cases, the accuracy of using $O$ as the order parameter is comparable to that of the conventional magnetization $M$ for the same lattice sizes. We note that a similar data collapse procedure has also been applied to $\delta O$ to extract the critical exponents $\gamma$, which produces consistent results that agree well with known exact values. More details can be found in Appendix B.
III.2 $q=3$ AFM Potts model on the diced lattice
@@
Figure 4: The diced lattice, as shown in (a), consists of a honeycomb sublattice (red circles, majority sites) and a triangular sublattice (blue circles, minority sites), where the latter is formed by sites located at the centers of the hexagons. Each blue circle is connected to six neighboring red circles, forming a bipartite structure. In order to map the spin configurations on the diced lattice into matrix form, an oblique rectilinear coordinate frame is used. In this coordinate frame, a blue circle appears after every two consecutive red circles along either the $x$ or $y$ direction. In (b) a snapshot of the spin configuration on $L=18$ diced lattice for the $q=3$ AFM Potts model is shown. It is taken from the MC sampling after the equilibration at the temperature $T=1.0\times 10^{-5}$. Three colors (red, blue and yellow) represent three spin values. A FM long-range order of one spin component (blue) is clearly observed on the triangular sublattice. In contrast, the spins on the honeycomb sublattice appear to randomly adopt the other two values (red and yellow).
@@
Figure 5: Same quantities as in Fig. 2, but for the $q=3$ AFM Potts model on the diced lattice. The vertical dashed lines in (a) and (b) mark the position of the critical temperature $T_{c}$ obtained in Ref. [12].
@@
The critical temperature $T_{c}$ and the critical exponents $\nu$ and $\beta$, determined by the data collapse method, are shown in Fig. 5(c). The estimated critical temperature $T_{c}=0.507$ is in excellent agreement with the reference values. The critical exponents, $\nu=0.77$, with an uncertainty range of $[0.75,0.79]$ and $\beta=0.110$, are also consistent with those of the $q=3$ FM Potts model ($\nu=5/6\approx 0.833$, $\beta=1/9\approx 0.111$). These observations corroborate the conclusion in Ref. [12] that the critical behavior of the model belongs to the universality class of the $q=3$ Potts ferromagnet. The data collapse of $\delta O$, along with the extracted critical exponents $\nu$ and $\gamma$ on this model, is presented in Appendix B.